# SuperCarver: Texture-Consistent 3D Geometry Super-Resolution for High-Fidelity Surface Detail Generation

### v1

@@ Abstract
Traditional production workflow of high-precision 3D mesh assets necessitates a cumbersome and laborious process of manual sculpting by specialized modelers. The recent years have witnessed remarkable advances in AI-empowered 3D content creation. However, although the latest state-of-the-arts are already capable of generating plausible structures and intricate appearances from images or text prompts, the actual mesh surfaces are typically over-smoothing and lack geometric details. This paper introduces SuperCarver, a 3D geometry super-resolution framework particularly tailored for adding texture-consistent surface details to given coarse meshes. Technically, we start by rendering the original textured mesh into the image domain from multiple viewpoints. To achieve geometric detail generation, we develop a deterministic prior-guided normal diffusion model fine-tuned on a carefully curated dataset of paired low-poly and high-poly normal renderings. To optimize mesh structures from potentially imperfect normal map predictions, we design a simple yet effective noise-resistant inverse rendering scheme based on distance field deformation. Extensive experiments show that SuperCarver generates realistic and expressive surface details as depicted by specific texture appearances, making it a powerful tool for automatically upgrading massive outdated low-quality assets and shortening the iteration cycle of high-quality mesh production in practical applications.
Figure 1: Given a 3D textured model lacking geometric details as input, SuperCarver generates realistic and expressive 3D surface details consistent with its texture appearance within 1 $\sim$ 2 minutes, significantly enhancing the production efficiency of high-quality mesh assets. (Row 1: The input 3D textured meshes. Rows 2&4: The original coarse surfaces and normal maps. Rows 3&5: The resulting high-fidelity mesh surfaces and normal maps after our geometry super-resolution.)
footnotetext: Corresponding author.
1 Introduction
3D modeling plays a key role in industrial design and digital entertainment fields, such as movies, video games, robotics, and virtual/augmented reality. In practice, the accuracy and quality of 3D mesh models directly influence specific downstream performances. However, creating high-quality mesh assets depends heavily on the manual efforts of specialized 3D modelers, which can be cumbersome and laborious.
Generally, the complete industrial 3D modeling pipeline can be broadly divided into three sequential stages in terms of the geometric complexity of the corresponding mesh surface: 1) Given initial concept designs, e.g., 2D sketch drawings, a mid-poly mesh is established to roughly replicate the overall spatial structure. 2) Subsequently, surface sculpting is performed to produce highly expressive surface details on a high-poly mesh, which typically consists of million-scale faces. 3) To ensure the efficiency of the rendering engine, a low-poly mesh with significantly simplified faces is created based on the high-poly mesh, followed by UV unwrapping, detail (such as normal/bump map) baking, and texture painting. Among the above procedures, the effects of high-poly geometry sculpting greatly determine the ultimate graphics quality and visual realism, but it accounts for a large portion of the entire manual workload. Therefore, there is an urgent need to particularly investigate the automated generation of high-fidelity surface details.
The recent advancements of deep learning-based generative modeling have led to remarkable strides in 3D content generation [51]. Restricted by insufficient 3D training data, early zero-shot architectures such as DreamFields [37] and DreamFusion [78] resort to distilling prior knowledge from pre-trained vision-language multi-modal models [84] or 2D image diffusion models [87]. However, such per-shape optimization process can typically suffer from inefficiency and inter-view inconsistency problems. Since the emergence of large-scale mesh datasets [23, 22], feed-forward 3D generation frameworks such as LRM [35] and InstantMesh [117] have received increasing attention, which can achieve end-to-end training and finish inference within several seconds. To further enhance generation quality without significantly compromising efficiency, another family of research such as SyncDreamer [58], Wonder3D [59], and Unique3D [113] proposes to apply multi-view diffusion models to synthesize 2D novel views and then reconstruct the target 3D mesh via iterative optimization procedures. In addition, analogous to 2D image diffusion, native 3D latent diffusion architectures such as CLAY [127] also demonstrate tremendous potential. Overall, the current state-of-the-art 3D generation methods are already capable of producing plausible geometric structures and intricate texture appearances. However, the actual mesh surfaces tend to be over-smoothing and lack geometric details, as shown in Figure 2. Moreover, it is worth noting that existing 3D generation approaches are designed for creating new 3D assets from fundamental creative elements, i.e., reasoning the underlying 3D structures from 2D images or text prompts. This means that they cannot be directly applied on a given 3D model for further enhancement, because there is no guarantee that the original structural characteristics could still be well preserved. In addition to image/text-to-3D generation paradigms, there also exist several lines of relevant studies, such as point cloud upsampling [123, 81], reference-based [33, 10] or text-guided [85, 29] 3D editing, geometric (e.g., depth/normal) map [2, 41, 27] or PBR material [104, 107] estimation. However, there is still a lack of specialized explorations for enhancing the surface details of specifically given 3D models.
In this paper, we present SuperCarver, a novel geometry super-resolution framework operating on a given 3D model for generating realistic and expressive surface details as depicted by its texture appearance. Considering the difficulties of directly learning 3D structural details, we choose normal maps, a 2.5D representation format, to efficiently bridge 2D and 3D domains. Accordingly, we start by rendering the input textured mesh into image space, which facilitates leveraging the power and pre-trained knowledge of image diffusion architectures. Still, accurately and robustly predicting normal maps through diffusion faces several challenges. On the one hand, normals are a form of sensitive and vulnerable geometric property, i.e., moderate numerical variations can lead to obvious damage to the underlying surface. Besides, without sufficient priors, normals predicted in image space often fail to align with the actual surface structures of given meshes. On the other hand, due to the inherent stochasticity of diffusion models, it is inevitable that normal predictions vary across different inferences. To address these problems, inspired by previous attempts to stabilize [45, 116] and accelerate [124, 108] image diffusion mechanisms, we adopt a chain of deterministic interpolations from rendered coarse normal maps to desired high-precision normal maps. In the meantime, necessary conditioning signals are incorporated by extracting semantic and geometric priors from rendered color images and depth maps.
After enhancing normal maps from multiple viewpoints, it is straightforward to perform gradient-based mesh updating [43]. Since optimizing from potentially imperfect multi-view predicted normal maps is prone to causing severe surface degradation, we customize a simple yet effective noise-resistant inverse rendering approach based on distance field deformation to deduce the corresponding surface details.
We conduct comprehensive performance evaluations and comparisons on a manually selected subset of textured mesh models from Objaverse [23] and ABO [21] datasets, which strongly demonstrate the effectiveness of our approach. We also verify that SuperCarver can be applied on off-the-shelf image/text-to-3D frameworks to further enhance the geometry quality of their generation results.
Figure 2: The generation results of the current state-of-the-arts, (a) InstantMesh [117] & (b) CLAY [127], typically lack high-fidelity surface details. The corresponding geometry enhancement results of our proposed SuperCarver are shown in (c) & (d).
2 Related Work
2.1 3D Content Generation
3D content generation [93, 51, 54] has always been a highly valuable yet rather challenging task. Unfortunately, the limited data volume and diversity of classic 3D shape repositories [115, 8] largely restrict the advancement of this field. In order to bypass the reliance on large-scale 3D training data, zero-shot schemes [37, 69, 71] are designed to constrain the text-prompted 3D generation process using the off-the-shelf language-image prior of CLIP [84]. With the advent of pre-trained 2D image diffusion models [34, 95, 87], score distillation [78, 105, 68, 53] is investigated to inject text guidance into the optimization procedure of various neural field representations [70, 42]. Subsequent researches further make continuous progress in improving generation quality and/or speed [11, 110, 62, 19, 100, 120, 49, 83], enabling 3D editing [85, 97], and supplementing image conditioning capabilities [66, 98, 80, 92]. Nevertheless, the distillation-based generation paradigm necessitates sophisticated processing to deal with inefficiency and inter-view inconsistency.
Thanks to the significant increase in the scale of 3D data [23, 22], it becomes promising to directly train feed-forward 3D generation networks. The pioneering work of LRM [35] builds an end-to-end regression framework to predict a triplane [7] NeRF representation, while Instant3D [46] develops a two-stage framework that begins with generating the “view-compiled” image grid from texts. Numerous follow-up studies continue to improve performances from different perspectives [103], such as shifting to the 3DGS [42] representation (LGM [99]), replacing transformers with convolutional architectures (CRM [111]), and adding more geometric supervisions (InstantMesh [117]).
To further boost the generation quality while maintaining satisfactory efficiency, viewpoint-conditioned 2D diffusion [56, 55] is explored to synthesize novel 2D views, followed by iterative optimization for sparse-view 3D reconstruction. However, lifting imperfect multi-view 2D predictions to 3D can be non-trivial and usually problematic. To ensure multi-view consistency, SyncDreamer [58] designs synchronized multi-view noise prediction by constructing an intermediate 3D feature volume. In Zero123++ [91], multiple images are tiled into a single frame to model the joint distribution of all pre-defined views. Wonder3D [59] builds a new multi-view cross-domain 2D diffusion pipeline controlled by a domain switcher to generate multi-view normal maps and color images, after which explicit 3D geometry is extracted via normal fusion. Era3D [47] incorporates explicit regression of camera parameters and designs more efficient and compact multi-view attention. Unique3D [113] constructs a powerful image-to-3D framework with comprehensive improvements in multi-view generation and upscaling of color images and normal maps, together with an efficient multi-stage mesh reconstruction algorithm.
Analogous to 2D diffusion for image generation, another promising direction is to develop native 3D diffusion architectures [44, 114, 125, 50, 127] for conditional 3D generation with different design choices of latent space representations. Additionally, auto-regressive generation of triangle sequences [94, 12, 112, 101, 15] and UV-domain diffusion [118, 26] driven by regular geometry image representations [30, 128, 129] also demonstrate new potential.
@@ 2 Related Work
2.2 Inverse Rendering and Mesh Optimization
@@
Inverse rendering aims at reasoning the underlying physical attributes of a target 3D scene from its 2D image observations. Different 3D scene parameters such as geometry, texture, material, and lighting can be optimized via backpropagating the gradients w.r.t. the 2D rendering outputs [72, 32]. Earlier research investigates various differentiable renderers [60, 40, 57, 13]. Later studies [5, 6, 96, 130, 126, 38] devote to learning neural field representations. For gradient-based mesh optimization, the most straightforward way is to explicitly perform surface deformation from pre-defined mesh templates [106, 31]. Nevertheless, directly optimizing the mesh vertex positions is prone to surface defect and degeneracy, thus requiring careful initialization, regularization, and remeshing [73, 76]. Another choice is to employ occupancy and distance fields [67, 13, 77] for implicit surface modeling. Since traditional iso-surface extraction algorithms [61, 25, 39, 74] cannot propagate gradients, various differentiable schemes [16, 18] are designed to operate on discrete occupancy grids (DMC [52]), continuous signed distance fields (MeshSDF [86]), or deformable tetrahedrons (DefTet [28], DMTet [89]). Moreover, NJF [1] depicts the mappings between pairs of source and target triangles via Jacobian matrices. FlexiCubes [90] incorporates carefully-chosen learnable parameters into [74] for achieving much more flexible local adjustment of mesh vertices.
2.3 3D Geometry Enhancement
@@
Similar to image super-resolution [109], point cloud upsampling [123, 121, 81, 48, 79, 82, 65, 129] can produce a denser set of spatial points uniformly located on the underlying 3D surface from sparse and possibly noisy inputs. However, these approaches are rarely capable of generating new geometric details. When a high-quality source mesh is given as the reference, detail transfer [4, 33] can synthesize similar geometric patterns on the target surface. More recent works [17, 9, 10] further apply such an examplar-based paradigm on coarse voxels for few-shot 3D shape detailization. Text-guided mesh editing frameworks [29, 3, 19] facilitate flexible and convenient geometry manipulation, but face challenges in accurate and fine-grained control. Additionally, PBR material generation [14, 122, 104, 107] typically involves estimating normal/bump maps, which can be potentially exploited for refinement purposes.
3 Proposed Method
Figure 3: The overall workflow of our proposed SuperCarver for texture-consistent geometry super-resolution. Given an input 3D textured model, we start by rendering color, depth, and normal maps, which are fed into a customized deterministic prior-guided diffusion architecture to generate detail-boosted normal maps. To perform gradient-based mesh refinement from potentially imperfect multi-view normal map predictions, we design a noise-resistant distance field deformation process to drive the optimization of high-fidelity 3D surface details.
3.1 Overall Workflow

### v2

@@ Abstract
Conventional production workflow of high-precision mesh assets necessitates a cumbersome and laborious process of manual sculpting by specialized 3D artists/modelers. The recent years have witnessed remarkable advances in AI-empowered 3D content creation for generating plausible structures and intricate appearances from images or text prompts. However, synthesizing realistic surface details still poses great challenges, and enhancing the geometry fidelity of existing lower-quality 3D meshes (instead of image/text-to-3D generation) remains an open problem. In this paper, we introduce SuperCarver, a 3D geometry super-resolution pipeline for supplementing texture-consistent surface details onto a given coarse mesh. We start by rendering the original textured mesh into the image domain from multiple viewpoints. To achieve detail boosting, we construct a deterministic prior-guided normal diffusion model, which is fine-tuned on a carefully curated dataset of paired detail-lacking and detail-rich normal map renderings. To update mesh surfaces from potentially imperfect normal map predictions, we design a noise-resistant inverse rendering scheme through deformable distance field. Experiments demonstrate that our SuperCarver is capable of generating realistic and expressive surface details depicted by the actual texture appearance, making it a powerful tool to both upgrade historical low-quality 3D assets and reduce the workload of sculpting high-poly meshes.
Index Terms:
@@
3D geometry super-resolution, diffusion model, mesh optimization, inverse rendering.
I Introduction
3D modeling plays a critical role in various industrial design and digital entertainment fields, such as movies, animations, video games, robotics, and virtual/augmented reality. In practice, the geometric accuracy of mesh assets directly influence the expressiveness of visual presentation and the performance of specific downstream tasks. Nevertheless, the production of high-quality mesh models heavily depends on manual efforts of specialized 3D modelers, which can be highly cumbersome and labor-intensive.
Generally, the complete industrial 3D modeling pipeline can be broadly divided into three sequential stages in terms of the geometric complexity of mesh surfaces:
  • 1)
    Mid-Poly: This stage primarily aims at converting initial concept design works (typically 2D sketch drawings) to 3D meshes, roughly reproducing the designer’s vision in terms of the overall style and spatial structure.
  • 2)
    High-Poly: To produce high-quality geometric surfaces, the preceding mid-poly meshes need to be carved within specialized digital sculpting software (typically ZBrush) to obtain high-poly meshes with realistic and expressive surface details.
  • 3)
    Low-Poly: To satisfy efficiency requirements of varying tasks and devices, high-poly meshes need to be significantly simplified to newly create low-poly meshes with typically thousands or hundreds of polygons, followed by UV unwrapping, detail (normal/bump map) baking, and texture painting.
Among the above three modeling stages, the effects of high-poly surface sculpting dominate the ultimate graphics quality and visual realism, while accounting for the majority of the overall workload. Therefore, there is an urgent need to develop automated surface detail generation techniques to accelerate the production of high-fidelity mesh production.
The recent advancements of deep learning-based generative modeling have led to remarkable strides in 3D content generation [1]. Early zero-shot architectures such as DreamFields [2] and DreamFusion [3] distill prior knowledge from pre-trained vision-language multi-modal models [4] and/or 2D image diffusion models [5]. Nevertheless, such per-shape optimization process can typically suffer from inefficiency and inter-view inconsistency problems. Since the emergence of large-scale mesh datasets [6, 7], feed-forward 3D generation frameworks such as LRM [8] and InstantMesh [9] receive increasing attention, which achieve end-to-end training and finish inference within several seconds. To further enhance generation quality without significantly compromising efficiency, another family of research such as SyncDreamer [10], Wonder3D [11], and Unique3D [12] proposes to apply multi-view diffusion models to synthesize 2D novel views and then reconstruct the target 3D mesh through iterative optimization procedures. Moreover, native 3D latent diffusion architectures such as CLAY [13] also demonstrate tremendous potential. The latest state-of-the-arts [14, 15, 16, 17] have continuously advanced the geometric accuracy of the generated meshes to new heights.
Functionally, previous 3D generation frameworks are particularly designed for creating new 3D assets according to low-dimensional textual (1D-to-3D) or visual (2D-to-3D) prompts. However, there still exists a large amount of practical demands to upgrade the geometric fidelity of existing 3D assets while strictly preserving the original structural characteristics, which is a different mesh-to-mesh (3D-to-3D) task scenario.
Additionally, there also exist several lines of relevant studies, such as point cloud upsampling [18, 19], reference-based [20] and text-guided [21] 3D editing, geometric (depth/normal) map [22, 23, 24] and PBR material [25] estimation, but there is still an obvious lack of specialized explorations for 3D surface detail generation.
Figure 1: Given an input 3D textured model lacking geometric details, SuperCarver generates realistic and expressive 3D surface details consistent with its texture appearance within 1 $\sim$ 2 minutes, significantly enhancing the production efficiency of high-quality mesh assets. (Row 1: The input 3D textured meshes. Rows 2&4: The original coarse surfaces and normal maps. Rows 3&5: The resulting high-fidelity mesh surfaces and normal maps after our geometry super-resolution.)
In this paper, we make the first attempt to introduce SuperCarver, a two-stage processing framework for geometry super-resolution operating on given 3D textured meshes to generate realistic and expressive surface details that are faithful to the corresponding texture appearances, as shown in Figure 1.
Considering the difficulties of directly learning 3D structural details, we choose normal maps, a 2.5D representation format, to efficiently bridge 2D and 3D domains. Accordingly, we start by rendering the input textured mesh into image space, which facilitates leveraging the power and pre-trained knowledge of 2D image diffusion architectures. Still, accurately and robustly predicting normal maps through diffusion faces several challenges, as outlined below:
  • 1)
    Normals are a form of sensitive and vulnerable geometric attributes. Moderate numerical perturbations can lead to obvious distortion and damage to the underlying surface. Without sufficient priors, normals predicted in the image space often fail to align with the actual surface structures of the given meshes.
  • 2)
    Due to the inherent stochasticity of diffusion models, it is inevitable that normal predictions exhibit certain degrees of jittering across multiple times of inferences.
To address these problems, inspired by previous attempts to stabilize [26] and accelerate [27] image diffusion mechanisms, we adopt a chain of deterministic interpolations from rendered detail-lacking and detail-rich normal maps. In the meantime, necessary conditioning signals are incorporated by extracting both semantic and geometric priors from rendered color images and depth maps.
Having enhanced the detail richness of normal maps from multiple viewpoints, it is straightforward to perform gradient-based mesh updating through the optimization loop of inverse rendering. However, optimizing the mesh surface from potentially imperfect multi-view normal map predictions is prone to causing unruly surface degradation. Therefore, we customize a simple yet effective noise-resistant inverse rendering approach based on distance field deformation to deduce the corresponding surface details.
Through comprehensive qualitative illustrations and quantitative evaluations, we demonstrate that our SuperCarver framework can supplement highly-expressive and texture-consistent surface details to various textured meshes with unsatisfactory geometric quality. The main contributions and superiorities of our work can be summarized as follows:
  • 1)
    We introduce the first mesh-to-mesh 3D geometry super-resolution learning framework. Our targeted task differs from previous 1D/2D-to-3D generative paradigms, offering new capabilities of upgrading the geometric quality of given mesh models in a non-destructive and identity-preserving manner.
  • 2)
    We propose a two-stage normal-centric pipeline featured by deterministic diffusion mechanism and noise-resistant distance field deformation.
  • 3)
    Our approach fills the gap in the current 3D generative modeling community for high-poly surface sculpting and historical low-quality mesh asset upgrading.
The remainder of this paper is organized as follows. Section II systematically reviews different perspectives of research on 3D generation, optimization, and enhancement. Section III introduces our SuperCarver framework for pixel-space normal boosting and geometry-space mesh updating. Section IV provides comprehensive evaluations, comparisons, and ablation studies to demonstrate the effectiveness of our approach. In the end, we conclude the whole paper in Section V.
II Related Work
II-A 3D Content Generation
3D content generation [28, 29, 30] has always been a highly valuable yet rather challenging task. Unfortunately, the limited data volume and the diversity of classic 3D shape repositories [31] largely restrict the advancement of this field. In order to bypass the reliance on large-scale 3D training data, zero-shot schemes [2, 32] are designed to constrain the text-prompted 3D generation process using the off-the-shelf language-image prior of CLIP [4]. With the advent of pre-trained 2D image diffusion models [5], score distillation [3] is investigated to inject text guidance [33] into the optimization procedure of various neural field representations [34, 35]. Subsequent researches further make continuous advancements in improving generation quality and/or speed [36, 37], enabling 3D editing [21, 38], and supplementing image conditioning capabilities [39, 40]. Nevertheless, the distillation-based generation paradigm necessitates sophisticated processing to deal with inefficiency and inter-view inconsistency.
Thanks to the significant increase in the scale of 3D data [6, 7], it becomes promising to directly train feed-forward 3D generation networks. The pioneering approach of LRM [8] constructs an end-to-end regression framework to predict a triplane [41] NeRF representation, while Instant3D [42] develops a two-stage framework that begins with generating the “view-compiled” image grid from texts. Numerous follow-up studies continue to improve performances from different perspectives [43, 44], such as shifting to strong 3DGS [35] representations (LGM [45]), replacing transformers with convolutional architectures (CRM [46]), and adding more geometric supervisions (InstantMesh [9]).
To further enhance the generation quality while maintaining satisfactory efficiency, viewpoint-conditioned 2D diffusion [47] is investigated to synthesize novel 2D views, followed by iterative optimization for sparse-view 3D reconstruction. However, lifting imperfect multi-view 2D predictions to 3D can be non-trivial and usually problematic. To ensure multi-view consistency, SyncDreamer [10] designs synchronized multi-view noise prediction by constructing an intermediate 3D feature volume. In Zero123++ [48], multiple images are tiled into a single 2D frame to model the joint distribution of all the pre-defined views. Wonder3D [11] builds an effective multi-view cross-domain 2D diffusion pipeline controlled by a domain switcher to generate multi-view normal maps and color images, after which explicit 3D geometry is extracted via normal fusion. Era3D [49] incorporates explicit regression of camera parameters and designs more efficient and compact multi-view attention. Unique3D [12] constructs a powerful image-to-3D framework with comprehensive improvements in multi-view generation and upscaling of color images and normal maps, together with an efficient multi-stage surface reconstruction algorithm.
Analogous to 2D diffusion for image generation, another promising direction is to develop native 3D diffusion architectures [50, 13] for conditional 3D generation with different design choices of latent space representations. Additionally, auto-regressive generation of triangle sequences [51, 52] and UV-domain diffusion [53] driven by regular geometry image representations [54, 55, 56] also demonstrate new potential.
II-B Inverse Rendering and Mesh Optimization
@@
Inverse rendering aims at reasoning the underlying physical attributes of a target 3D scene from its 2D image observations. Different 3D scene parameters such as geometry, texture, material, and lighting can be optimized via backpropagating the gradients w.r.t. the 2D rendering outputs [57, 58]. Earlier research investigates various differentiable renderers [59, 60]. Later studies [61, 62, 63] devote to learning neural field representations. For gradient-based mesh optimization, the most straightforward way is to explicitly perform surface deformation from pre-defined mesh templates [64, 65]. Nevertheless, directly optimizing the mesh vertex positions is prone to surface defect and degeneracy, thus requiring careful initialization, regularization, and remeshing [66, 67, 68]. Another choice is to employ occupancy and distance fields [69, 70, 71] for implicit surface modeling. Since traditional iso-surface extraction algorithms [72, 73] cannot propagate gradients, various differentiable schemes [74, 75] are designed to operate on discrete occupancy grids (DMC [76]), continuous signed distance fields (MeshSDF [77]), or deformable tetrahedrons (DMTet [78]). Moreover, NJF [79] depicts the mappings between pairs of source and target triangles via Jacobian matrices. FlexiCubes [80] incorporates carefully-chosen learnable parameters into [73] for achieving much more flexible local adjustment of mesh vertices.
Figure 2: The overall processing pipeline of our proposed SuperCarver for texture-consistent geometry super-resolution. Given an input 3D textured model, we begin by rendering color, depth, and normal maps, which are fed into a customized deterministic prior-guided diffusion architecture to generate detail-boosted normal maps. To perform gradient-based mesh refinement from potentially imperfect multi-view normal map predictions, we further design a noise-resistant distance field deformation process to drive the optimization of high-fidelity 3D surface details.
II-C 3D Geometry Enhancement
@@
Similar to image super-resolution, 3D point cloud upsampling [18, 19, 81, 82] can produce a denser set of spatial points uniformly located on the underlying 3D surface from sparse and possibly noisy inputs. However, these approaches are rarely capable of generating new geometric details. When a high-quality source mesh is given as the reference, detail transfer [83] can synthesize similar geometric patterns on the target surface. More recent works [20] further apply such an examplar-based paradigm on coarse voxels for few-shot 3D shape detailization. Text-guided mesh editing frameworks [84] facilitate flexible and convenient geometry manipulation, but face challenges in accurate and fine-grained control. Additionally, PBR material generation [25] typically involves estimating normal/bump maps, which can be potentially exploited for refinement purposes.
III Proposed Method
III-A Overall Workflow
For a 3D textured mesh model $\mathcal{M}$ with vertices $\mathcal{V}$ and faces $\mathcal{F}$, we aim at generating high-fidelity surface details reflected by its specific texture appearance, thus producing a high-poly version $\mathcal{M}^{u}=(\mathcal{V}^{u},\mathcal{F}^{u})$ of the input coarse mesh surface $\mathcal{M}$. As a geometry sculpting process, it is particularly required that the original spatial structures should be preserved.